A)Write the chemical name and symbol corresponding to each of the following theoretical descriptions :
1.30 electrons and a net charge of 2 +

B)Write out the formula of the oxide that must be dissolved in water to make each of the following acids or bases :

1.aluminum hydroxide
2.sulfuric acid
Answer #1

0 Great Atomic number = 32 Number of electory - Afonuc neember-2 = 32-2 = 30 +3720 - 2Al(OH)2 Aluminium hydroxide (Al2O3 Alum